# Fine-grained Spatial-temporal MLP Architecture for Metro Origin-Destination Prediction

### v1

@@
Abstract.
Accurate prediction of metro traffic is crucial for optimizing metro scheduling and enhancing overall transport efficiency. Analyzing fine-grained and comprehensive relations among stations effectively is imperative for metro Origin-Destination (OD) prediction. However, existing metro OD models either mix information from multiple OD pairs from the station’s perspective or exclusively focus on a subset of OD pairs. These approaches may overlook fine-grained relations among OD pairs, leading to difficulties in predicting potential anomalous conditions. To address these challenges, we analyze traffic variations from the perspective of all OD pairs and propose a fine-grained spatial-temporal MLP architecture for metro OD prediction, namely ODMixer. Specifically, our ODMixer has double-branch structure and involves the Channel Mixer, the Multi-view Mixer, and the Bidirectional Trend Learner. The Channel Mixer aims to capture short-term temporal relations among OD pairs, the Multi-view Mixer concentrates on capturing relations from both origin and destination perspectives. To model long-term temporal relations, we introduce the Bidirectional Trend Learner. Extensive experiments on two large-scale metro OD prediction datasets HZMOD and SHMO demonstrate the advantages of our ODMixer. The code will be available.
Origin-Destination Prediction, Spatial-temporal Learning, Metro System
copyright: noneccs: Applied computing Transportationccs: Applied computing Forecastingccs: Information systems Spatial-temporal systemsccs: Information systems Decision support systems
1. Introduction
@@
Figure 1. Comparison of different views and models for metro OD prediction. (a) illustrates the difference in encoding the OD Matrix from the station view and the OD pair view. N is the number of stations, T is the number of time intervals. (b) shows various models for processing the $N^{2}$ tokens.
@@
With the rapid development of urbanization, city populations are surging, accompanied by a gradual increase in the number of vehicles (Jin et al., 2023; Liu et al., 2022b; Wang et al., 2023; Lin et al., 2023; Li et al., 2024). To enhance daily commuting convenience, people are increasingly considering the metro as their preferred mode of transportation (Liu et al., 2023; Zhu et al., 2022). Simultaneously, cities are proactively expanding their metro networks, aiming to augment the city’s transport capacity and alleviate traffic pressures, thereby enhancing the quality of life for urban residents. However, to ensure the operational efficiency of the metro system and the seamless movement of passengers, real-time monitoring of metro passenger flow and predicting future trends have become imperative to address various emergencies. Consequently, metro OD forecasting, directed at estimating the flow among all stations in the forthcoming period, emerges as a critical task in metro scheduling and flow management.
@@
Over time, various models utilizing deep learning have been employed to address the metro OD prediction problem (Liu et al., 2022c; Gong et al., 2020; Xu et al., 2023; Shen et al., 2024). However, as illustrated in Fig. 1(a), these models predominantly adopt a station view, encoding all data from each station at every time moment in a coarse-grained manner. This approach results in the mixing of data from various OD pairs, causing the encoded features to encompass information from different OD pairs. Consequently, these models face challenges in discerning the flow changes of OD pairs. In reality, given that the purpose of most people’s trips is clear, and the origin and destination of the majority of trips are predetermined, it is more suitable to encode the OD matrix from the perspective of OD pairs. The OD pair view encodes each OD pair independently to distinguish different OD pairs, which is more fine-grained than the station view that mixes multiple OD pairs.
@@
Actually, there are exist some models that predict the metro OD matrix from the OD pair view (Zhang et al., 2023b; Yang et al., 2022), but they only focus a subset of OD pairs in the metro. This constrained approach neglects the richness of information among all OD pairs, resulting in the model’s inability to learn fine-grained relations across the entirety of OD pairs. Consequently, it hampers the model’s capability to address potential emergencies, comprehend the dynamics of traffic changes, making it difficult to provide effective support for the management of the metro. To achieve more comprehensive metro OD prediction, it is imperative to consider information from all OD pairs. However, considering the number of OD pairs as $N^{2}$, the time complexity needed to calcaulate the relations among $N^{2}$ OD pairs using Transformer or to construct a virtual graph among OD pairs and subsequently employ Graph Convolution Network (GCN) for information aggregation is $O(N^{4})$, as shown in Fig. 1(b). Consequently, keeping the balance between model’s parameters and computation time while considering all OD pairs remains challenging.
@@
To tackle the aforementioned challenges, we propose a simple yet efficient fine-grained spatial-temporal MLP model named ODMixer. Initially, we individually encode the time series of OD pairs base on OD pair view. Subsequently, to capture the intricate relations among OD pairs, we introduce two modules, Channel Mixer and Multi-view Mixer. The Channel Mixer module facilitates the modeling of temporal correlation by calculating relations among features of OD pairs in the temporal dimension. Considering the substantial time required for direct computation of relations among all OD pairs, as with methods like Transformer, and the potential confusion and difficulty in distinguishing important associations, the Multi-view Mixer calculates relations among different OD pairs from two perspectives—Origin and Destination. This enhances the comprehensive understanding of OD pair relations, thereby improving the efficiency of modeling fine-grained relations among OD pairs. Additionally, to exploit the time-periodicity of metro traffic, we introduce the Bidirectional Trend Learner that effectively perceives the long-term temporal relations between traffic flow and enhance the temporal attributes of features.
Our contributions can be summarized as follows:
@@
2. Related Work
2.1. Origin-Destination Prediction
OD prediction (Wang et al., 2024) involves the accurate estimation of traffic flow between two regions over a given period of time, presenting a formidable challenge due to its intricate and dynamic nature. As a pivotal aspect of urban traffic management, OD prediction has garnered extensive research attention.
As a special type of OD prediction, the metro OD prediction (Gong et al., 2020; Zhang et al., 2021; Liu et al., 2022c; Ye et al., 2023; Xu et al., 2023; Zhang et al., 2023b; Zhu et al., 2023) presents unique challenges compared to the general OD prediction, primarily because the precise destinations of passengers are not accurately known until they reach their destinations. This leads to incomplete data at the current moment, making it challenging to effectively depict the real-time traffic distribution. To address this issue and obtain more comprehensive data on traffic distribution, several studies focused on predicting unfinished orders to supplement the incomplete OD matrix. For instance, in (Liu et al., 2022c), long and short historical distributions were utilized to estimate the distribution of unfinished orders at the current time. This study integrated the completed and unfinished distributions in the feature space. In (Gong et al., 2020), an indication matrix was created, with values assigned based on the assumption that the travel time of each OD pair at each timestamp follows a normal distribution. However, this approach overlooked the unfinished OD pair data.
@@ 2.1. Origin-Destination Prediction
2.2. MLP-based Model
With the popularity of the Transformer, its exceptional performance has led to its widespread adoption across various fields (Liu et al., 2021b; Chen et al., 2023b; Tang et al., 2023; Wei et al., 2023; Yan et al., 2023; Liu et al., 2022a, 2021c, 2019, 2018a, 2018b; Xu et al., 2022). However, several MLP-based models have attracted attention for addressing diverse problems recently. In computer vision, (Tolstikhin et al., 2021) addressed vision challenges through an entirely MLP-based implementation. Another notable model gMLP (Liu et al., 2021a) outperformed some Transformer-based models by leveraging channel projection, spatial projection, and gating mechanisms. (Touvron et al., 2022) introduced a residual structure atop MLP, incorporating training methods like self-supervision for efficient training. (Tang et al., 2022) focused on reducing computational costs while maintaining performance. For sequential recommendation problems, (Li et al., 2022) employed a tri-directional mixing MLP model, while (Zhang et al., 2023a) introduced SpatialMixer and TemporalMixer to address spatio-temporal dependencies and temporal variations in traffic prediction. For time series forecasting task, (Ekambaram et al., 2023) and (Chen et al., 2023a) proposed MLP-based models dedicated to efficient time series forecasting.
Figure 2. (a) illustrates the architecture of double-branch ODMixer, both branches learn features using ODIM. Subsequently, the two branches interact with each other using the BTL, contributing to the final output. (b) depicts the BTL module, which aims to jointly model the features from both branches, thereby enhancing the information exchange. (c) represents the ODIM module, which models the temporal attributes of the OD pair features using the Channel Mixer, while the comprehensive relations among OD pairs are learned using the Multi-view Mixer. (d) shows the Channel Mixer, Origin Mixer, Des Mixer.
@@ 2.2. MLP-based Model
3. Preliminaries
In this section, a concise introduction is provided for the notation and problem definitions. Table 1 presents a comprehensive list of commonly used symbols for reference.
@@ 3. Preliminaries
(1) $\mathrm{OD}_{t+1}^{d}=f(\{\mathrm{OD}^{d-1},\mathrm{IOD}^{d},\mathrm{unf}^{d}\}_{t-T+i})$
@@ 4. Methodology
Table 1. Notation Table.
@@ 4.1. Architecture Overview
4.2. OD Matrix Processing
In this subsection, we introduce the approach for estimating the current UOD distribution by leveraging the historical UOD distribution and the fine-grained encoding of the OD matrix from the OD pair perspective, supported by the Embedding Layer.
4.2.1. Unfinished Orders Processing
@@
(2) $\mathrm{\widetilde{UOD}}=\frac{\mathrm{unf}\ast(\mathrm{UOD}_{S}+\mathrm{UOD}_{L})}{2}$
@@
(3) $\mathrm{OD}=\mathrm{IOD}+\mathrm{\widetilde{UOD}}$
@@ 4.2.2. Embedding Layer
(4) $\boldsymbol{H_{i}}=\boldsymbol{W^{E}}\mathrm{OD}_{i}$
@@
4.3. Origin-Destination Interact Module (ODIM)
@@ 4.3.1. Channel Mixer
(5) $\boldsymbol{H^{C}_{i}}=LayerNorm(\boldsymbol{H_{i}}+\boldsymbol{W^{C_{2}}}\sigma(\boldsymbol{W^{C_{1}}}\boldsymbol{H_{i}}))$
where, $i=1,\cdots,N^{2}$, $\sigma$ is the activation function, $\boldsymbol{W^{C_{1}}}\in\mathbb{R}^{d_{C}\times d}$ is learnable weights of the first fully connected layer in the Channel Mixer, $\boldsymbol{W^{C_{2}}}\in\mathbb{R}^{d\times d_{C}}$ is learnable weights of the second fully connected layer, $\boldsymbol{H^{C}}\in\mathbb{R}^{N\times N\times d}$ represents the feature after temporal attributes have enhanced by the Channel Mixer. We employ layer normlization (LayerNorm) (Ba et al., 2016) and residual connection (He et al., 2016) in the Channel Mixer. The Channel Mixer structure is specifically designed to merge attributes from diverse time intervals within the feature representation, enhancing the model’s ability to capture temporal traffic patterns. Most importantly, the Channel Mixer parameters are shared across all OD pairs, implicitly fostering temporal interactions between distinct OD pairs without necessitating an increase in parameter count. This shared parameter design contributes to the model’s efficiency in learning global temporal patterns, thereby enhancing its capacity to comprehend and predict variations in traffic over time.
4.3.2. Multi-view Mixer
@@
(6) $\boldsymbol{H^{O}_{i}}$ $=\boldsymbol{W^{O_{2}}}\sigma(\boldsymbol{W^{O_{1}}}\boldsymbol{H^{C_{O}}_{i}})$
where $i=1,\cdots,N\times d$, $\boldsymbol{W^{O_{1}}}\in\mathbb{R}^{d_{O}\times d}$ and $\boldsymbol{W^{O_{2}}}\in\mathbb{R}^{d\times d_{O}}$ are the learnable weights of the Origin Mixer, $H_{O}\in\mathbb{R}^{N\times N\times d}$ denotes the output of the module.
@@
(7) $\boldsymbol{H^{D}_{i}}$ $=\boldsymbol{W^{D_{2}}}\sigma(\boldsymbol{W^{D_{1}}}\boldsymbol{H^{C_{D}}_{i}})$
Where, $i=1,\cdots,N\times d$, $\boldsymbol{W^{D_{1}}}\in\mathbb{R}^{d_{D}\times d}$ and $\boldsymbol{W^{D_{2}}}\in\mathbb{R}^{d\times d_{D}}$ are the learnable weights of the Des Mixer, $\boldsymbol{H^{D}}\in\mathbb{R}^{N\times N\times d}$ is the output of the Des Mixer.
@@
(8) $\boldsymbol{H^{F}}=LayerNorm(\boldsymbol{H^{O}}+\boldsymbol{H^{D}}+\boldsymbol{H^{C}})$
@@
4.4. Bidirectional Trend Learner (BTL)

### v2

# ODMixer: Fine-grained Spatial-temporal MLP for Metro Origin-Destination Prediction

@@
Abstract
Metro Origin-Destination (OD) prediction is a crucial yet challenging spatial-temporal prediction task in urban computing, which aims to accurately forecast cross-station ridership for optimizing metro scheduling and enhancing overall transport efficiency. Analyzing fine-grained and comprehensive relations among stations effectively is imperative for metro OD prediction. However, existing metro OD models either mix information from multiple OD pairs from the station’s perspective or exclusively focus on a subset of OD pairs. These approaches may overlook fine-grained relations among OD pairs, leading to difficulties in predicting potential anomalous conditions. To address these challenges, we analyze traffic variations from the perspective of all OD pairs and propose a fine-grained spatial-temporal MLP architecture for metro OD prediction, namely ODMixer. Specifically, our ODMixer has double-branch structure and involves the Channel Mixer, the Multi-view Mixer, and the Bidirectional Trend Learner. The Channel Mixer aims to capture short-term temporal relations among OD pairs, the Multi-view Mixer concentrates on capturing relations from both origin and destination perspectives. To model long-term temporal relations, we introduce the Bidirectional Trend Learner. Extensive experiments on two large-scale metro OD prediction datasets HZMOD and SHMO demonstrate the advantages of our ODMixer. Our code is available at https://github.com/KLatitude/ODMixer.
Index Terms:
Correlation Learning, Spatial-temporal Learning, Origin-Destination Prediction, Metro System.
I Introduction
@@
Figure 1: Comparison of different views and models for metro OD prediction. (a) illustrates the difference in encoding the OD Matrix from the station view and the OD pair view. N is the number of stations, T is the number of time intervals. (b) shows various models for processing the $N^{2}$ tokens.
@@
With the rapid development of urbanization, city populations are surging, leading to a gradual increase in the number of vehicles [1, 2, 3, 4]. To enhance daily commuting convenience, metro is increasingly considered as the preferred transportation mode. Concurrently, cities are proactively expanding their metro networks to augment transport capacity and alleviate traffic pressures, thereby enhancing the quality of life for urban residents. In some metropolises, such as Beijing and Shanghai, nearly 10 million metro travel transactions occur daily. This immense ridership poses significant challenges for metro operations. Consequently, accurately forecasting future ridership is crucial for effective metro scheduling and route planning. However, to ensure the operational efficiency of the metro system and the seamless movement of passengers, real-time monitoring of metro passenger flow and predicting future trends have become imperative to address various emergencies. Consequently, metro OD prediction, directed at estimating the flow among all stations in the forthcoming period, emerges as a critical task in metro scheduling and flow management.
Compared to general traffic OD prediction, metro OD prediction has unique challenges. First, due to varying time intervals between passengers entering and exiting stations, it is impossible to count the orders of passengers who have not yet exited when calculating the OD matrix at any given time, resulting in an incomplete OD matrix. This incompleteness is a significant characteristic of metro OD prediction. Secondly, the large number of metro stations leads to a high-dimensional OD matrix, posing substantial challenges to model design. Additionally, the uneven distribution of metro traffic causes significant differences between OD pairs, resulting in a sparse OD matrix. Analyzing these traffic differences is challenging.
(a) Flow difference between OD pairs
(b) Interval difference between OD pairs
Figure 2: Comparison of flow changes between OD pairs with the same origin. The x-axis represents July 4 and 5, 2019, the y-axis represents the flow value between OD pairs. The OD:(190,196) in the legend means the origin is station 190 and the destination is station 196. Figure (a) shows that although the OD pairs with the same origin have similar change trends, the flow values are very different. Figure (b) illustrates that the flow rate trends of two OD pairs with the origin exhibit significant differences over time.
@@
Over time, various deep learning models have been employed to address the metro OD prediction problem [5, 6, 7, 8]. However, as illustrated in Fig. 1(a), these models predominantly adopt a station view, encoding all data from each station at every time moment in a coarse-grained manner. This approach results in the mixing of data from various OD pairs, causing the encoded features to encompass information from different OD pairs. Consequently, these models face challenges in discerning the flow changes of OD pairs. As shown in Figure 2(a), although the time variation trends of the flow rates of two OD pairs with the origin are similar, the difference in flow rates is very pronounced. Additionally, as shown in Figure 2(b), the time variation trends between the two OD pairs are completely different. Therefore, if the OD matrix is modeled based on the perspective of the stations, the differences between different OD pairs may be ignored, making the model unable to effectively capture the temporal changes in flow, resulting in a decrease in model performance. In reality, given that the purpose of most people’s trips is clear, and the origin and destination of the majority of trips are predetermined, it is more suitable to encode the OD matrix from the perspective of OD pairs. The OD pair view encodes each OD pair independently to distinguish different OD pairs, which is more fine-grained than the station view that mixes multiple OD pairs.
@@
Actually, there exists some models that predict the metro OD matrix from the OD pair view [9, 10], but they only focus a subset of OD pairs in the metro. This constrained approach neglects the richness of information among all OD pairs, resulting in the model’s inability to learn fine-grained relations across the entirety of OD pairs. Consequently, it hampers the model’s capability to address potential emergencies, comprehend the dynamics of traffic changes, making it difficult to provide effective support for the management of the metro. To achieve more comprehensive metro OD prediction, it is imperative to consider information from all OD pairs. However, considering the number of OD pairs as $N^{2}$, the time complexity needed to calculate the relations among $N^{2}$ OD pairs using Transformer or to construct a virtual graph among OD pairs and subsequently employ Graph Convolution Network (GCN) for information aggregation is $O(N^{4})$, as shown in Fig. 1(b). Consequently, keeping the balance between model’s parameters and computation time while considering all OD pairs remains challenging.
@@
Inspired by the efficiency of the MLP architecture [11, 12], we propose a simple yet efficient fine-grained spatial-temporal MLP model named ODMixer, to tackle the aforementioned challenges. Initially, we individually encode the time series of OD pairs base on OD pair view. Subsequently, to capture the intricate relations among OD pairs, we introduce two modules, Channel Mixer and Multi-view Mixer. The Channel Mixer module facilitates the modeling of temporal correlation by calculating relations among features of OD pairs in the temporal dimension. Considering the substantial time required for direct computation of relations among all OD pairs, as with methods like Transformer, and the potential confusion and difficulty in distinguishing important associations, the Multi-view Mixer calculates relations among different OD pairs from two perspectives: Origin and Destination. This enhances the comprehensive understanding of OD pair relations, thereby improving the efficiency of modeling fine-grained relations among OD pairs. Additionally, to exploit the temporal periodicity of metro traffic, we introduce the Bidirectional Trend Learner that perceives the long-term temporal relations between traffic flow and enhance the temporal attributes of features. Our contributions are summarized as follows:
@@
II Related Work
II-A Traffic Flow Prediction
Traffic flow prediction is a very important task in smart city systems. Its main task is to use historical traffic information to accurately predict future traffic flow. There are many methods to solve this problem. Initially, researchers used traditional statistical methods to analyze traffic changes, such as ARIMA[13]. Later, some machine learning-based methods began to emerge, such as SVM[14] and decision trees[15].
With the rapid development of deep learning, neural networks have been used to solve traffic prediction problems. These representative works include many structural types, mainly including CNN [16, 17], RNN [18, 19], Attention [20, 21], etc. DeepSD[22] proposed an end-to-end deep learning framework and integrates external environment information to realize the automatic discovery of complex supply-demand patterns from the car-hailing service data. Based on three temporal relations, MDL[23] used two three-stream fully convolutional networks (3S-FCNs) to achieve flow prediction for points and edges. DeepCrowd[24] divided the area into multiple fine-grained grids, and combines ConvLSTM and attention mechanisms to analyze and fuse flow information at multiple times to achieve Large-Scale Citywide Crowd Density and Flow Prediction.
Later, the GCN was popular to process the relation between nodes because of the natural network structure in the traffic flow prediction problem [25, 26, 27]. STDGRL[28] designed a spatio-temporal dynamic graph relation learning model. The spatial relation is captured dynamically by dynamically learning the graph structure between nodes, and the long-term temporal relation is learned by using attention. STWave+[29] utilized decoupling technology to partition one-hour traffic data into stable trends and fluctuating events. This approach employs a dual-channel spatiotemporal network to separately model these components, leveraging self-supervised learning and contrastive loss to propagate long-term trend information to hourly trends. Ultimately, it forecasts forthcoming traffic conditions through adaptive fusion techniques.
However, these methods focus solely on the total flow at each node, they fail to analyze the flow relations between nodes. This limitation restricts the information that can be obtained, making it less suitable for practical applications such as traffic analysis and control. To address this issue, we concentrate on the relations between nodes. Compared to focusing only on the total flow at each node, this OD relation provides a more fine-grained analysis. Additionally, the total flow at a node can be derived from the OD flow, highlighting the greater value of OD flow data.
II-B Origin-Destination Prediction
OD prediction [30] involves the accurate estimation of traffic flow between two regions over a given period of time, presenting a formidable challenge due to its intricate and dynamic nature. As a pivotal aspect of urban traffic management, OD prediction has garnered extensive attention.
As a special type of OD prediction, the metro OD prediction [6, 31, 5, 32, 7, 9, 33] presents unique challenges compared to the general OD prediction, primarily because the precise destinations of passengers are not accurately known until they reach their destinations. This leads to incomplete data at the current moment, making it challenging to effectively depict the real-time traffic distribution. To address this issue and obtain more comprehensive data on traffic distribution, several studies focused on predicting unfinished orders to supplement the incomplete OD matrix. For instance, in HIAM[5], long and short historical distributions were utilized to estimate the distribution of unfinished orders at the current time. This study integrated the completed and unfinished distributions in the feature space. In Gong et al.[6], an indication matrix was created, with values assigned based on the assumption that the travel time of each OD pair at each timestamp follows a normal distribution. However, this approach overlooked the unfinished OD pair data.
@@ II-B Origin-Destination Prediction
II-C MLP-based Model
With the popularity of the Transformer, its exceptional performance has led to its widespread adoption across various fields. However, several MLP-based models have attracted attention for addressing diverse problems recently. In computer vision, MLP-Mixer[11] addressed vision challenges through an entirely MLP-based implementation. Another notable model gMLP [34] outperformed some Transformer-based models by leveraging channel projection, spatial projection, and gating mechanisms. ResMLP[35] introduced a residual structure atop MLP, incorporating training methods like self-supervision for efficient training. sMLPNet[36] focused on reducing computational costs while maintaining performance. For sequential recommendation problems, MLP4Rec[37] employed a tri-directional mixing MLP model, while MLPST[38] introduced SpatialMixer and TemporalMixer to address spatio-temporal dependencies and temporal variations in traffic prediction. For time series forecasting task, Ekambaram et al.[39] and Chen et al.[40] proposed MLP-based models dedicated to efficient time series forecasting.
Figure 3: (a) illustrates the architecture of double-branch ODMixer, both branches learn features using ODIM. Subsequently, the two branches interact with each other using the BTL, contributing to the final output. (b) depicts the BTL module, which aims to jointly model the features from both branches, thereby enhancing the information exchange. (c) represents the ODIM module, which models the temporal attributes of the OD pair features using the Channel Mixer, while the comprehensive relations among OD pairs are learned using the Multi-view Mixer. (d) shows the Channel Mixer, Origin Mixer, Des Mixer.
@@ II-C MLP-based Model
III Preliminaries
In this section, a concise introduction is provided for the notation and problem definitions. Table I presents a comprehensive list of commonly used symbols for reference.
@@ III Preliminaries
$\mathrm{OD}_{t+1}^{d}=f(\{\mathrm{OD}^{d-1},\mathrm{IOD}^{d},\mathrm{unf}^{d}\}_{t-T+i})$ (1)
@@ IV Methodology
TABLE I: Notation Table.
@@ IV-A Architecture Overview
IV-B OD Matrix Processing
Since ODMixer contains two branches (prev and cur), the input of the prev branch is the OD matrix at the same time yesterday. Since yesterday’s order information has been completed, its input is a complete OD matrix and no operation is required. The input of the cur branch is the OD matrix at the current time, so it is incomplete and needs to be processed. In this subsection, we introduce the approach for estimating the current UOD distribution by leveraging the historical UOD distribution and the fine-grained encoding of the OD matrix from the OD pair perspective, using the Embedding Layer.
IV-B1 Unfinished Orders Processing
@@
$\mathrm{\widetilde{UOD}}=\frac{\mathrm{unf}\ast(\mathrm{UOD}_{S}+\mathrm{UOD}_{L})}{2}$ (2)
@@
$\mathrm{OD}=\mathrm{IOD}+\mathrm{\widetilde{UOD}}$ (3)
@@ IV-B2 Embedding Layer
$\boldsymbol{H_{i}}=\boldsymbol{W^{E}}\mathrm{OD}_{i}$ (4)
@@
IV-C Origin-Destination Interact Module (ODIM)
@@ IV-C1 Channel Mixer
$\boldsymbol{H^{C}_{i}}=LayerNorm(\boldsymbol{H_{i}}+\boldsymbol{W^{C_{2}}}\sigma(\boldsymbol{W^{C_{1}}}\boldsymbol{H_{i}}))$ (5)
where, $i=1,\cdots,N^{2}$, $\sigma$ is the activation function, $\boldsymbol{W^{C_{1}}}\in\mathbb{R}^{d_{C}\times d}$ is learnable weights of the first fully connected layer in the Channel Mixer, $\boldsymbol{W^{C_{2}}}\in\mathbb{R}^{d\times d_{C}}$ is learnable weights of the second fully connected layer, $\boldsymbol{H^{C}}\in\mathbb{R}^{N\times N\times d}$ represents the feature after temporal attributes have enhanced by the Channel Mixer. We employ layer normlization (LayerNorm) [41] and residual connection [42] in the Channel Mixer. The Channel Mixer structure is specifically designed to merge attributes from diverse time intervals within the feature representation, enhancing the model’s ability to capture temporal traffic patterns. Most importantly, the Channel Mixer parameters are shared across all OD pairs, implicitly fostering temporal interactions between distinct OD pairs without necessitating an increase in parameter count. This shared parameter design contributes to the model’s efficiency in learning global temporal patterns, thereby enhancing its capacity to comprehend and predict variations in traffic over time.
IV-C2 Multi-view Mixer
@@
$\boldsymbol{H^{O}_{i}}$ $=\boldsymbol{W^{O_{2}}}\sigma(\boldsymbol{W^{O_{1}}}\boldsymbol{H^{C_{O}}_{i}})$ (6)
where $i=1,\cdots,N\times d$, $\boldsymbol{W^{O_{1}}}\in\mathbb{R}^{d_{O}\times N}$ and $\boldsymbol{W^{O_{2}}}\in\mathbb{R}^{N\times d_{O}}$ are the learnable weights of the Origin Mixer, $H_{O}\in\mathbb{R}^{N\times N\times d}$ denotes the output of the module.
@@
$\boldsymbol{H^{D}_{i}}$ $=\boldsymbol{W^{D_{2}}}\sigma(\boldsymbol{W^{D_{1}}}\boldsymbol{H^{C_{D}}_{i}})$ (7)
Where, $i=1,\cdots,N\times d$, $\boldsymbol{W^{D_{1}}}\in\mathbb{R}^{d_{D}\times N}$ and $\boldsymbol{W^{D_{2}}}\in\mathbb{R}^{N\times d_{D}}$ are the learnable weights of the Des Mixer, $\boldsymbol{H^{D}}\in\mathbb{R}^{N\times N\times d}$ is the output of the Des Mixer.
@@
$\boldsymbol{H^{F}}=LayerNorm(\boldsymbol{H^{O}}+\boldsymbol{H^{D}}+\boldsymbol{H^{C}})$ (8)
@@
IV-D Bidirectional Trend Learner (BTL)